Currently, the two fastest trains are the Japanese Maglev and the French TGV. The sum of their fastest speeds is 718.2 mph. IF the speed of the Maglev is 3.8mph faster thean the speed of TGV, find the speeds of each.
Answer by checkley77(12844)

X+X+3.8=718.2
2X=718.2-3.8
2X=714.4
2X=714.4
X=714.4/2
X=357.2 MPH FOR THE TGV TRAIN.
357.2+3.8=361 MPH. FOR THE MAGLEV TRAIN.